# thanks: These two authors contributed equally to this workthanks: These two authors contributed equally to this work

### v1

@@ III Results: Overview
Figure 2: Overview of the dimensionless ratio $\xi/L$ for different $\sigma$ values ($\xi$ refers to the $2^{\text{nd}}$-moment correlation length). The phase transition type changes from a second-order transition to a BKT-type transition for $\sigma>2$, as shown here. The $2^{\text{nd}}$-moment correlation length divided by $L$, $\xi/L$, is plotted as a function of inverse temperature $\beta$ for various $\sigma$ values: (a) $\sigma=1.25$, (b) $\sigma=1.75$, (c) $\sigma=1.875$, (d) $\sigma=2$, (e) $\sigma=2.1$, and (f) $\sigma=3$. System sizes range from $L=64$ to $L=4096$. For $\sigma<2$, the curves for different system sizes intersect, indicating a second-order phase transition. In the $\sigma=2.1$ case, smaller system sizes (up to $L=1024$) show a crossing due to the strong finite-size effect. However, for larger system sizes ($L=2048$ and $L=4096$), the $\xi/L$ values tend to overlap beyond the crossing point, consistent with a BKT transition. For $\sigma=3$, all curves for $L>64$ overlap after the transition point, confirming the BKT transition.
@@
Before delving into the detailed analysis of the LR XY model, we provide an overview of the main findings to facilitate a better understanding. As the long-range model enters the non-classical regime in $\sigma\leq\sigma_{*}$, the critical behavior differs from that of the short-range regime. For the LR Ising model, the universal properties of the transition, such as the critical exponents and other universal quantities, vary as $\sigma$ in the non-classical regime. Therefore, previous studies of the LR Ising model focus on the accurate extraction of critical exponents or the scaling behavior of universal values, which, however, can be extremely subtle [8, 11, 10]. In contrast, for the LR XY model, we expect a much more evident signature when the system enters the non-classical regime — the type of transition changes from a BKT transition to a second-order transition. The change in transition type leads to drastically different behaviors in both the low-temperature and high-temperature phases and the critical points, making determining $\sigma_{*}$ much easier than the LR Ising case.
@@
Based on extensive MC simulations from our previous work [19] and the present study, we have identified three distinct regimes in the phase diagram of the LR XY model: the classical regime ($\sigma\leq 1$), the non-classical regime ($1<\sigma\leq 2$), and the short-range regime ($\sigma>2$), as illustrated in Fig.1. For $\sigma<1$, the system undergoes a second-order transition to a LRO low-$T$ phase, with critical behavior described by Gaussian mean-field theory [46]. In the short-range regime with $\sigma>2$, the system exhibits a standard BKT transition to a QLRO phase. The non-classical regime ($1<\sigma\leq 2$) is particularly noteworthy, where detailed FSS analysis reveals a second-order phase transition to a ferromagnetic phase at low-$T$ with $\sigma$-dependent critical exponents depending. The phase diagram obtained in this work is inconsistent with the previously suggested one from field theoretical analysis — the proposed intermediate QLRO phase in the regime $1.75<\sigma\leq 2$ is not observed from our data [17, 18]. Moreover, our results suggest the crossover point between SR and LR universality is at $\sigma=2$, instead of the $\sigma=1.75$ given by Sak’s criterion.
@@
The second-order phase transition in the non-classical regime is evident from the correlation length ratio $\xi/L$ as a function of inverse temperature $\beta$ for various $\sigma$. Figure 2 presents $\xi/L$ versus $\beta$ for $\sigma=1.25$, $1.75$, $1.875$, $2$, $2.1$ and $3$. The data from system sizes ranging from $L=64$ to $L=4096$ are shown for a comprehensive examination of finite-size effects. For $\sigma\leq 2$, $\xi/L$ curves of different $L$s clearly exhibit a universal intersection at some finite temperature, below which $\xi/L$ diverges as the system size increases. This typical scaling behavior indicates a second-order phase transition into an LRO phase, and the crossing points mark the location of the critical points $T_{c}$. This differs from the BKT transition in the SR regime, for which the results are also shown as a comparison. For $\sigma=3$, the $\xi/L$ curves do not show a crossing; instead, curves for system sizes greater than $L=64$ converge to a universal function at low-$T$, confirming a BKT transition to a QLRO phase. However, at $\sigma=2.1$, a more nuanced behavior occurs. While the $\xi/L$ curves of smaller system sizes (up to $L=1024$) seemly suggest a crossing, while those for larger sizes ($L=2048$ and $L=4096$) deviate from this trend and overlap after the critical point. Such behavior reflected the strong finite-size effect near the boundary between LR and SR universality. Nevertheless, as demonstrated in later sections, the low-$T$ and high-$T$ properties of the model further confirm that the system undergoes a BKT transition for $\sigma=2.1$.
IV Results I: Long-range order and Goldstone mode at low temperature
@@
In this section, we conduct a comprehensive study of the system’s low-temperature properties. First, in Sec. IV.1, we establish the existence of LRO in the low-$T$ phase for $\sigma\leq 2$. Following this, in Sec. IV.2, we explore the Goldstone mode excitations in the low-$T$ phase for $\sigma<2$, demonstrating that the correlation function adheres to the form $g(r)=g_{0}+cr^{-\eta_{\ell}}$ where $\eta_{\ell}=2-\sigma$. For $\sigma=2$, we propose a possible logarithmic behavior and substantiate it numerically. Subsequently, the QLRO phase at low-$T$s for $\sigma>2$ is discussed in Sec IV.3. Finally, we revisit the boundary case of $\sigma=2$ and examine its distinctive finite-size scaling in Sec. IV.4. Our research primarily focuses on $\sigma$ values of $1.75,1.875,2$ at various low temperatures characterized by $\beta=1,2,4,8$. Additionally, cases of $\sigma=2.1,3$ are partially included for comparison.
@@ IV.1 Long-Range Order for $\sigma\leq 2$
Figure 3: Different behaviors of $(\xi/L)^{2}$ for various $\sigma$. $(\xi/L)^{2}$ is plotted in the semi-logarithmic coordinates as a function of system size $L$ is shown for various $\sigma$ values: $\sigma=1.75$ (green circle), $1.875$ (green upward triangle), $2$ (red square), $2.1$ (blue downward triangle), $2.2$ (blue diamond), and $3$ (blue pentagon), at inverse temperatures $\beta=1$ (a) and $\beta=2$ (b). A power-law and logarithmic dependence of $(\xi/L)^{2}$ on $L$ is observed for $\sigma<2$ and $\sigma=2$, indicating a divergence as the system size increases, which reflects LRO in the system, whereas, for $\sigma>2$, $(\xi/L)^{2}$ converges to a constant, indicating QLRO in the system.
@@
In this part, we demonstrate the existence of LRO in the low-$T$ phase for $\sigma\leq 2$. We first study the scaling behavior of the squared correlation length ratio $(\xi/L)^{2}$, which exhibits distinct behavior in the LRO and QLRO phases. Specifically, for given $T<T_{c}$, as the system size increases, this ratio diverges in the LRO phase, while in a QLRO phase, it converges to some non-zero constant. Fig. 3 displays $(\xi/L)^{2}$ as a function of $L$ at inverse temperatures $\beta=1$ and $\beta=2$ in a semi-logarithmic scale, revealing the distinct low-$T$ scaling behaviors of $(\xi/L)^{2}$ for various values of $\sigma$. For $\sigma<2$, the curves bend upwards as $L$ increases, and this divergence can be well-fitted by a power-law function, which will be discussed later. For $\sigma=2$, the data points fall onto a straight line, suggesting that $(\xi/L)^{2}$ diverges logarithmically with $L$. For both cases, despite the different forms of scaling, $(\xi/L)^{2}$ diverges as $L$ increases, indicating the presence of LRO in the low-$T$ phase for $\sigma\leq 2$. In contrast, for $\sigma>2$, the system is in the QLRO phase at low temperatures, and $(\xi/L)^{2}$ are expected to converge. This can be seen clearly in the case of $\sigma=3$ at both $\beta=1$ and $2$ where $(\xi/L)^{2}$ rapidly converges to a constant as $L$ increases. However, for $\sigma=2.1$ and $2.2$, the proximity to the marginal case $\sigma=2$ results in strong finite-size effects, leading to a slow convergence of $(\xi/L)^{2}$. As shown in Fig 3, for both $\beta=1$ and $2$, $(\xi/L)^{2}$ curves slowly increase at a slackening rate, and we expect them to finally converge to a constant. In Sec. IV.3, we further verify this expectation.
@@
Figure 4: The existence of spontaneous magnetization for $\sigma\leq 2$ in low-$T$ phase. The plots of $\langle M^{2}\rangle$ as a function of system size $L$ for various inverse temperatures $\beta=1$ (red), $2$ (blue), $4$ (green), and $8$ (black), at $\sigma=1.75$ (a1), $\sigma=1.875$ (b1), and $\sigma=2$ (c1). The $x$-axis representations vary by panel: $L^{-\eta_{\ell}}$ ($\eta_{\ell}=0.25,0.125$ for $\sigma=1.75,1.875$, respectively) in panels (a1) and (b1), $1/\ln(L/L_{0})$ in panel (c1). The fitting parameter $L_{0}$ equals $e^{-5.4},e^{-6.8},e^{-6.38},e^{-6.2}$ for $\beta=1,2,4,8$ in panel (c1). As $L$ increases, the system approaches the thermodynamic limit, indicating spontaneous magnetization at low temperatures. The figure also includes corresponding plots for $M_{r}^{2}$ in (a2) through (c2) where $M^{2}_{r}$ are defined as $M_{r}^{2}=\langle M^{2}\rangle-b\langle M_{k}^{2}\rangle$. The $x$-axis representations vary by panel: $L^{-0.65}$ in panel (a2), $L^{-0.5}$ in panel (b2), and $L^{-0.4}$ in panel (c2). $b$ equals $22$ for panel (a2), 48 for panel (b2), $149,175,152,154$ in $\beta=1,2,4,8$ respectively for panel (c2). After subtracting a positive quantity, $M_{r}^{2}$ continues to reach a positive value at the thermodynamic limit with smaller finite-size corrections, which is strong evidence of spontaneous magnetization.
@@
The finite-size scaling of $\langle M^{2}\rangle$ provides another evidence of the LRO in the low-$T$ phase. In the LRO phase, $\langle M^{2}\rangle$ converges to a positive value in the thermodynamic limit, while in the QLRO phase, it reduces to 00 as the system size increases. To extrapolate the value of $\langle M^{2}\rangle$ in the thermodynamic limit, we fit the data at finite system sizes. For the LR XY model, due to the spontaneous breaking of continuous symmetry, the LRO phase shall exhibit Goldstone modes. It is shown in Sec IV.2 that the correlation function has the form: $g(x)=g_{0}+c\cdot x^{-\eta_{\ell}}$ where $\eta_{\ell}=2-\sigma$. Therefore, considering the relationship between $\langle M^{2}\rangle$ and correlation function: $\langle M^{2}\rangle\sim\frac{1}{L^{2}}\int g(r)\mathrm{d}^{2}r$, we fit the $\langle M^{2}\rangle$ with following ansatz,
@@
where $a_{1}L^{-\omega}$ term accounts for additional corrections. For $\sigma=2$, however, the correlation function can’t be easily obtained from a Goldstone mode analysis due to the degeneracy between the LR and the SR interaction terms; nevertheless, in Sec. IV.2, we propose a possible scaling form at $\sigma=2$: $\chi_{k}\sim L^{2}\ln(L/L_{0})$ or $\langle M_{k}^{2}\rangle\sim\ln(L/L_{0})$. Taking into account that $M_{k}$ is the Fourier mode of $M$, $\langle M^{2}\rangle$ is likely to exhibit scaling behaviors similar to $\langle M_{k}^{2}\rangle$, while its constant term remains preserved. Thus for $\sigma=2$, $\langle M^{2}\rangle$ should behave as
@@
First three panels of Fig. 4 display the extrapolation of $M^{2}$ to the $L\rightarrow\infty$ limit, at various temperatures for $\sigma=1.75$ (panel a1), $\sigma=1.875$ (panel b1), and $\sigma=2$ (panel c1). According to Eq. (19), the leading correction term of $\langle M^{2}\rangle$ is $L^{-\eta_{\ell}}$, thus in panel (a1) and (b1) ($\sigma=1.75$ and 1.875), we plot the $x$-axis as $L^{-\eta_{\ell}}$. The plots reveal an asymptotic linear relationship, except for data of small system size, which deviates due to finite-size effects. The non-zero intercepts in the $y$-axis reveal the presence of spontaneous magnetization in the low-$T$ phases. Panel (c1) illustrates the logarithmic scaling $\langle M^{2}\rangle$ for $\sigma=2$, where we plot $\langle M^{2}\rangle$ versus $1/\ln(L/L_{0})$, where $L_{0}$ is a fitting parameter. The data fit well to the ansatz in Eq. (20), and spontaneous magnetization is evident from the non-zero intercept. Moreover, the linear relationship observed in this panel confirms the logarithmic scaling of $M^{2}$ as described by Eq. (20).
Extrapolating $\langle M^{2}\rangle$ to the thermodynamic limit in the first three panels in Fig. 4 demonstrates the existence of LRO in the low-$T$ phase for $\sigma\leq 2$. However, despite the large simulated system size, the slowly decaying corrections result in notable discrepancies between the data and the extrapolated results, particularly in the case of $\sigma=2$. To mitigate these finite-size corrections, we define the residual magnetization,
@@
where $a_{1}L^{-\omega}$ attributes to correction terms. A notable difference between the scaling behavior between $\langle M^{2}\rangle$ and $\langle M_{k}^{2}\rangle$ is that $\langle M_{k}^{2}\rangle$ vanishes in the thermodynamic limit. Similarly for $\sigma=2$, we expect $\langle M_{k}^{2}\rangle$ behaves as,
$\langle M_{k}^{2}\rangle=\frac{a_{0}}{\ln(L/L_{0})},$ (23)
again, a detailed discussion on this scaling behavior is provided in the next subsection IV.2. Since $\langle M_{k}^{2}\rangle$ shares the same scaling as the leading correction term in $\langle M^{2}\rangle$, selecting an appropriate value for $b$ can effectively cancel out the leading correction, thus reducing the finite-size effects in $M_{r}^{2}$. More importantly, because $\langle M_{k}^{2}\rangle$ dwindles and finally vanishes in the thermodynamic limit, for finite system size, $M^{2}_{r}$ is a lower bound of $\langle M^{2}\rangle$, and in the $L\rightarrow\infty$ limit, one has $M_{r}^{2}=\langle M^{2}\rangle$.
In practice, the fitting of equations, such as Eq. (19) and Eq. (23), provides the value of $a_{0}$, which represents the amplitude of the correction term we aim to eliminate. The ratio of $a_{0}$ in $\langle M^{2}\rangle$ and $\langle M_{k}^{2}\rangle$ yields the value of $b$. It is important to note that $b$ remains almost constant for various low temperatures, ensuring this method is broadly applicable in the low-$T$ phase. The residual magnetization $M_{r}^{2}$ is then fitted with the ansatz,
@@
The lower three panels of Fig. 4 display $M_{r}^{2}$ as a function of $L^{-\omega}$ for $\sigma=1.75,1.875,2$, with the values of $b$ and $\omega$ provided in the figure caption. Thanks to the reduced finite-size corrections, the data points are closer to the extrapolated thermodynamic limit than previous results, and the extrapolated curve clearly intersects the $y$-axis at positive values, indicating finite magnetization density. The discrepancies at smaller $L$ values indicate the presence of additional, smaller corrections beyond $L^{-\omega}$. Furthermore, unlike $\langle M^{2}\rangle$, the coefficient of the leading correction term of $M_{r}^{2}$ is negative, causing $M_{r}^{2}$ to increase with the size of the system. In all cases, the values of $M_{r}^{2}$ for the largest system size are already positive. Consequently, in the $L\rightarrow\infty$ limit, one must have $M^{2}\geq M_{r}^{2}>0$. Thus, we conclude that the system retains long-range order in the low-$T$ phase for $\sigma\leq 2$.

### v2

@@ III Overview of Results
Figure 2: Overview of the dimensionless ratio $\xi/L$ for different $\sigma$ values. The phase transition type changes from a second-order transition for $\sigma\leq 2$ to a BKT-type transition for $\sigma>2$, as shown here. The second-moment correlation length divided by $L$, $\xi/L$, is plotted as a function of inverse temperature $\beta$ for various $\sigma$ values: (a) $\sigma=1.25$, (b) $\sigma=1.75$, (c) $\sigma=1.875$, (d) $\sigma=2$, (e) $\sigma=2.1$, and (f) $\sigma=3$. System sizes range from $L=64$ to $L=4096$. For $\sigma\leq 2$, the curves for different system sizes exhibits a clear intersection, indicating a second-order phase transition. For $\sigma=3$, all $\xi/L$ curves for $L>64$ converge after the transition point, a hallmark of a BKT transition. In the $\sigma=2.1$ case, smaller system sizes (up to $L=1024$) show a crossing due to the strong finite-size effect. For larger system sizes ($L=2048$ and $L=4096$), the $\xi/L$ values tend to converge in the low-T phase. This subtle case will be elucidated in Sec. IV.3.
@@
Before delving into the detailed analysis of the LR XY model, we provide an overview of the main findings to facilitate a better understanding. As the long-range model enters the nonclassical regime in $\sigma\leq\sigma_{*}$, the critical behavior differs from that of the short-range regime. For the LR Ising model, the universal properties of the transition, such as the critical exponents and other universal quantities, vary as $\sigma$ in the nonclassical regime. Therefore, previous studies of the LR Ising model focus on the accurate extraction of critical exponents or values of universal ratios like Binder cumulant, which, however, can be extremely subtle and challenging [6, 9, 8]. In contrast, for the LR XY model, we expect a much more evident signature when the system enters the nonclassical regime — the type of transition changes from a BKT transition to a second-order transition. The change in transition type leads to qualitatively distinct behaviors in both the low- and high-T phases, as well as in the critical region. As a result, determining the crossover point $\sigma_{*}$ becomes considerably easier in the LR XY model than in the LR Ising case.
@@
Based on extensive MC simulations, we identify three distinct regimes in the phase diagram of the LR XY model: the classical regime ($\sigma\leq 1$), the nonclassical regime ($1<\sigma\leq 2$), and the short-range regime ($\sigma>2$), as illustrated in Fig.1. For $\sigma<1$, the system undergoes a second-order transition to a LRO low-T phase, with critical behavior governed by Gaussian mean-field theory [49]. In the short-range regime with $\sigma>2$, the system exhibits a standard BKT transition to a QLRO phase. The nonclassical regime ($1<\sigma\leq 2$) is particularly noteworthy, where detailed FSS analysis reveals a second-order phase transition to a ferromagnetic phase at low T with $\sigma$-dependent critical exponents depending. This finding contrasts with the previously suggested phase diagram based on field theoretical analysis — the proposed intermediate QLRO phase in the regime $1.75<\sigma\leq 2$ is not observed from our data [19, 20, 50]. Moreover, our results suggest the crossover point between SR and LR universality is at $\sigma=2$, rather than at $\sigma=1.75$ as predicted by Sak’s criterion.
@@
The second-order phase transition in the nonclassical regime is strongly hinted in Fig. 2 where the correlation length ratio $\xi/L$ is plotted as a function of the inverse temperature $\beta$. Figure 2 presents $\xi/L$ versus $\beta$ for $\sigma=1.25$, $1.75$, $1.875$, $2$, $2.1$ and $3$. The data from system sizes ranging from $L=64$ to $L=4096$ are shown for a comprehensive examination of finite-size effects. For $\sigma\leq 2$, $\xi/L$ curves of different $L$s clearly exhibit a universal intersection at some finite temperature, and in the low-T phase $\xi/L$ diverges as the system size increases, reflecting the emergence of long-range order. This typical scaling behavior indicates a second-order phase transition into an LRO phase, and the crossing points mark the location of the critical points $T_{c}$. This differs from the BKT transition in the SR regime, for which the results are also shown as a comparison. For $\sigma=3$, the $\xi/L$ curves do not show a crossing; instead, curves for system sizes greater than $L=64$ converge to a universal function at low temperatures, confirming a BKT transition to a QLRO phase. However, at $\sigma=2.1$, the behavior is more nuanced. While the $\xi/L$ curves of smaller system sizes (up to $L=1024$) seemly suggest a crossing, those for larger sizes ($L=2048$ and $L=4096$) start to converge after the critical point. Such behavior reflected the strong finite-size effect near the boundary between LR and SR universality. A systematic and detailed analysis will be given later. We will show that, in the low-T phase, the squared correlation-length ratio, $(\xi/L)^{2}$, diverges as $\sim L^{2-\sigma}$ for $\sigma<2$ and logarithmically as $\sim\ln L$ for $\sigma=2$; in contrast, $\xi/L$ statures to some finite and $T$-dependent value for $\sigma=2.1$. In the high-T phase when approaching the critical point, the correlation length $\xi$ exhibits power-law divergence for $\sigma\leq 2$, indicative of a second-order phase transition; in contrast, $\xi$ diverges exponentially for $\sigma=2.1$ and $3$, indicative of a BKT phase transition. Moreover, the fitted critical exponents also separate $\sigma\leq 2$ from the SR regime.
IV Long-range order and Goldstone mode at low temperature
@@
In this section, we conduct a comprehensive study of the system’s low-T properties. In subsection IV.1, we show the existence of LRO in the low-T phase for $\sigma\leq 2$. In subsection IV.2, we investigate the Goldstone mode excitations in the low-T phase for $\sigma<2$, demonstrating that the correlation function adheres to the form $g(r)=g_{0}+cr^{-\eta_{\ell}}$ where $\eta_{\ell}=2-\sigma$. For $\sigma=2$, we propose a possible logarithmic behavior and verify it numerically. Subsequently, the QLRO phase at low temperatures for $\sigma>2$ is discussed in subsection IV.3. Finally, we revisit the boundary case of $\sigma=2$ and examine its distinctive finite-size scaling in subsection IV.4. Our research primarily focuses on $\sigma$ values of $1.75,1.875,2$ at various low temperatures characterized by $\beta=1,2,4,8$. Additionally, cases of $\sigma=2.1,3$ are partially included for comparison.
@@ IV.1 Long-Range Order for $\sigma\leq 2$
Figure 3: Different behaviors of $(\xi/L)^{2}$ in the low-T phase for various $\sigma$. $(\xi/L)^{2}$ as a function of system size $L$ in the semi-logarithmic coordinates is shown for various $\sigma$ values at inverse temperatures $\beta=1$ (a) and $\beta=2$ (b). The linear behavior of the red squares demonstrates the logarithmic divergence of $(\xi/L)^{2}$ for the case of $\sigma=2$. For $\sigma<2$ (the green dots in the figure), $(\xi/L)^{2}$ diverges faster, indicating a possible power-law divergence. This is confirmed in Sec. IV.2. The diverging $(\xi/L)^{2}$ is a signature of the LRO in the low-T phase for $\sigma\leq 2$ For $\sigma>2$ (the blue dots), $(\xi/L)^{2}$ tends to converge, indicating QLRO in the system.
@@
In this part, we demonstrate the existence of LRO in the low-T phase for $\sigma\leq 2$. We first study the scaling behavior of the squared correlation length ratio $(\xi/L)^{2}$, which exhibits distinct behavior in the LRO and QLRO phases. Specifically, for given $T<T_{c}$, as the system size increases, this ratio diverges in the LRO phase, while in a QLRO phase, it converges to some non-zero constant. Fig. 3 displays $(\xi/L)^{2}$ as a function of $L$ at inverse temperatures $\beta=1$ and $\beta=2$ in a semi-logarithmic scale, revealing the distinct low-T scaling behaviors of $(\xi/L)^{2}$ for various values of $\sigma$. For $\sigma<2$, the curves bend upwards as $L$ increases, and this divergence can be well-fitted by a power-law function, which will be discussed later. For $\sigma=2$, the data points fall onto a straight line, suggesting that $(\xi/L)^{2}$ diverges logarithmically with $L$. For both cases, despite the different forms of scaling, $(\xi/L)^{2}$ diverges as $L$ increases, indicating the presence of LRO in the low-T phase for $\sigma\leq 2$. In contrast, for $\sigma>2$, the system is in the QLRO phase at low temperatures, and $(\xi/L)^{2}$ is expected to converge. This can be seen clearly in the case of $\sigma=3$ at both $\beta=1$ and $2$ where $(\xi/L)^{2}$ rapidly converges to a constant as $L$ increases. However, for $\sigma=2.1$ and $2.2$, the proximity to the marginal case $\sigma=2$ results in strong finite-size effects, leading to a slow convergence of $(\xi/L)^{2}$. As shown in Fig 3, for both $\beta=1$ and $2$, $(\xi/L)^{2}$ curves slowly increase at a slackening rate, and we expect them to finally converge to a constant. In subsection IV.3, we further verify this expectation.
@@
Figure 4: The existence of spontaneous magnetization for $\sigma\leq 2$ in low-T phase. The squared magnetization $\langle M^{2}\rangle$ is plotted as a function of system size $L$ for various temperatures at $\sigma=1.75$ (a1), $\sigma=1.875$ (b1), and $\sigma=2$ (c1). The $x$-axis represents the leading finite-size correction term, which varies by panel: $L^{-\eta_{\ell}}$ ($\eta_{\ell}=0.25,0.125$ for $\sigma=1.75,1.875$, respectively) in panels (a1) and (b1), $1/\ln(L/L_{0})$ in panel (c1). The fitting parameter $L_{0}=e^{-5.4},e^{-6.8},e^{-6.38},e^{-6.2}$ for $\beta=1,2,4,8$ in panel (c1). For all three panels, the fitting lines align with the data point, and intersect with y-axis at positive values, indicating the spontaneous magnetization at low temperatures in the thermodynamic limit. To diminish the finite-size correction, the plots of $M_{r}^{2}=\langle M^{2}\rangle-b\langle M_{k}^{2}\rangle$ are presented at $\sigma=1.75$ (a2), $\sigma=1.875$ (b2) and $\sigma=2$ (c2). The $x$-axis represents the leading correction term: $L^{-0.65}$ in panel (a2), $L^{-0.5}$ in panel (b2), and $L^{-0.4}$ in panel (c2). $b$ equals $22$ for panel (a2), 48 for panel (b2), $149,175,152,154$ in $\beta=1,2,4,8$ respectively for panel (c2). After subtracting a positive quantity, $M_{r}^{2}$ converges faster and continues to reach a positive value at the thermodynamic limit with smaller finite-size corrections, which is strong evidence of spontaneous magnetization.
The Monte Carlo data of the squared magnetization $\langle M^{2}\rangle$ are presented in Fig. 4 (a1-c1) for various $\beta$s. The squared magnetization is plotted against its leading correction term (will be explained below), demonstrating the extrapolation of $\langle M^{2}\rangle$ to the $L\rightarrow\infty$ limit. For $\sigma=1.75,1.875$ and $2$, at $\beta\geq 1$, $\langle M^{2}\rangle$ tends to converge to nonzero values in the thermodynamic limit, which strongly indicates the existence of LRO in the low-T phase for $\sigma\leq 2$. It will be shown that $\langle M^{2}\rangle$ follows a power-law convergence for $\sigma<2$ and a logarithmic convergence for $\sigma=2$, and the corresponding fitting form of $\langle M^{2}\rangle$ is explained below.
@@
For the LR XY model, due to the spontaneous breaking of continuous symmetry, the LRO phase shall exhibit Goldstone modes. It is shown in Sec IV.2 that the correlation function has the form: $g(x)=g_{0}+c\cdot x^{-\eta_{\ell}}$ where $\eta_{\ell}=2-\sigma$. Therefore, considering the relationship between $\langle M^{2}\rangle$ and correlation function: $\langle M^{2}\rangle\sim\frac{1}{L^{2}}\int g(r)\mathrm{d}^{2}r$, $\langle M^{2}\rangle$ is fitted with the following ansatz,
@@
where $a_{1}L^{-\omega}$ term accounts for additional corrections. For $\sigma=2$, however, the correlation function can’t be easily obtained from a Goldstone mode analysis due to the degeneracy between the LR and the SR interaction terms; nevertheless, in subsection IV.2, we propose a possible scaling form at $\sigma=2$: $\chi_{k}\sim L^{2}\ln(L/L_{0})$ or $\langle M_{k}^{2}\rangle\sim\ln(L/L_{0})$. Since $M_{k}$ is the Fourier mode of $M$, $\langle M^{2}\rangle$ is expected to exhibit similar scaling behavior, with an additional constant term. Thus for $\sigma=2$, $\langle M^{2}\rangle$ is fitted as
@@
In Fig. 4 (a1) and (b1) ($\sigma=1.75$ and $1.875$), we plot the $x$-axis as $L^{-\eta_{\ell}}$, the leading correction term according to Eq. (18). The plots reveal an asymptotic linear relationship, except for the small system size data, which is due to finite-size effects. The non-zero intercepts in the $y$-axis reveal the presence of spontaneous magnetization in the low-T phases. Fig. 4 (c1) illustrates the logarithmic scaling $\langle M^{2}\rangle$ for $\sigma=2$, where we plot $\langle M^{2}\rangle$ versus $1/\ln(L/L_{0})$, with $L_{0}$ a fitting parameter. The apparent linear relationship confirms the logarithmic scaling of $\langle M^{2}\rangle$ and the data fit well to the ansatz in Eq. (19). Again, the spontaneous magnetization is evident from the non-zero intercept.
Extrapolating $\langle M^{2}\rangle$ to the thermodynamic limit in the first three panels in Fig. 4 demonstrates the existence of LRO in the low-T phase for $\sigma\leq 2$. However, due to the large finite-size effect, for $\sigma=1.875$ and $2$, $\langle M^{2}\rangle$ converges slowly. To mitigate the finite-size corrections and better observe the LRO, we define the residual squared magnetization,
@@
where $a_{1}L^{-\omega}$ attributes to correction terms. A notable difference between the scaling behavior of $\langle M^{2}\rangle$ and $\langle M_{k}^{2}\rangle$ is that $\langle M_{k}^{2}\rangle$ vanishes in the thermodynamic limit. Similarly for $\sigma=2$, we expect $\langle M_{k}^{2}\rangle$ behaves as,
$\langle M_{k}^{2}\rangle=\frac{a_{0}}{\ln(L/L_{0})}.$ (22)
A detailed discussion on this scaling behavior will be provided in the next subsection IV.2. Since $\langle M_{k}^{2}\rangle$ shares the same scaling as the leading correction term in $\langle M^{2}\rangle$, selecting an appropriate value for $b$ can effectively cancel out the leading correction, thus reducing the finite-size effects in $M_{r}^{2}$. More importantly, since $b$ is a positive constant, $M^{2}_{r}$ act as a lower bound of $\langle M^{2}\rangle$, and in the $L\rightarrow\infty$ limit, one has $M_{r}^{2}=\langle M^{2}\rangle$. Hence, if $M_{r}^{2}$ converges to a positive value in the thermodynamic limit, $\langle M^{2}\rangle$ must also be positive.
The lower three panels of Fig. 4 display $M_{r}^{2}$ as a function of $L^{-\omega}$ for $\sigma=1.75,1.875,2$, with the values of $b$ and $\omega$ provided in the figure caption. With reduced finite-size corrections, the data points converge quickly to the extrapolated thermodynamic-limit value. The extrapolated curve clearly intersects the $y$-axis at positive values, indicating finite magnetization density. The discrepancies at smaller $L$ values indicate the presence of additional, smaller corrections beyond $L^{-\omega}$. Furthermore, the coefficient of the leading correction term of $M_{r}^{2}$ is negative, causing $M_{r}^{2}$ to increase with the size of the system. In all cases, the values of $M_{r}^{2}$ for the largest system size are already positive. Consequently, in the $L\rightarrow\infty$ limit, one must have $\langle M^{2}\rangle\geq M_{r}^{2}>0$. Thus, we conclude that the system retains long-range order in the low-T phase for $\sigma\leq 2$.
The fitting detail of $M_{r}^{2}$ is shown below. Firstly, the fitting of Eq. (18) and Eq. (22), provides the the amplitude of the correction term we aim to eliminate, namely, $a_{0}$. The ratio of $a_{0}$ in $\langle M^{2}\rangle$ and $\langle M_{k}^{2}\rangle$ determines the value of $b$ in $M_{r}^{2}$. $M_{r}^{2}$ is then fitted with the ansatz,
@@
It is worth noting that our conclusion partially contradicts the theoretical arguments presented in Ref. [17], stating that the 2D LR XY and Heisenberg system cannot sustain LRO at $\sigma=2$. This discrepancy between the theoretical predictions and our numerical evidence underscores the marginal case at $\sigma=2$ as an open and intriguing problem.